# Bidirectional Consistency Models

### v2

@@ A.3.2 Inpainting
We describe our inpainting process in Algorithm 6. In our experiments, we set the initial noise scale $s=0.5$. Additionally, instead of inverting the image back to $T=80.0$, we empirically find inverting the image to $2.0$ already suffice for satisfactory inpainting outcomes, and hence, we use a 3-step inversion (i.e., $N=4$), where $t_{1}=\varepsilon=0.07$, $t_{2}=0.4$, $t_{3}=1.0$ and $t_{4}=2.0$.
@@
As for the results by CMs, we implement Algorithm 4 as described by Song et al. [17] in their work. To ensure a fair comparison, we opt for the improved model, iCT [18], over the original CMs, since iCT delivers superior generation performance. Nonetheless, due to the absence of officially released codes and checkpoints by the authors, we reproduce and train our own iCT model. Our reproduced iCT yields an FID score closely matching that reported by Song and Dhariwal [18] — our reproduced FID is $2.87$ compared to the reported $2.83$ — affirming the reliability of our results.
A.3.3 Blind Restoration of Compressed Image
We compress CIFAR-10 test set with JPEG and RECOMBINER [30] at $5.46$ bpp and $0.92$ bpp (bits per pixel), respectively, which result in similar PSNR (peak signal-to-noise ratio) with different artifacts. For the JPEG restoration, we invert the input image with the 2-step sampling to reach the noise scale of $2.0$, where $\varepsilon=t_{1}=0.07,t_{2}=0.7,t_{3}=2.0$. For the INR restoration, we invert the input image with the 3-step sampling to the noise scale of $80.0$, where $\varepsilon=t_{1}=0.07,t_{2}=1.5,t_{3}=6.0,t_{4}=80.0$. We then apply single-step denoising to both settings.
A.3.4 Defending Black-Box Adversarial Attack
We use 1-step inversion with $\varepsilon=t_{1}=0.11,t_{2}=80.0$ and 1-step generation. For the 4 models used for black-box adversarial attack and defense, we use SGD optimizer with momentum $0.9$, weight decay $5\times 10^{-4}$ and batch size $128$ to train each of them for $200$ epochs. The learning rate is initially set to $0.1$ and decays to $0.01$ at epoch $100$ and further decays to $0.001$ at epoch $150$. We hold a validation set of $1,000$ samples from the full training set to select the better checkpoint. As a result, each of them achieves a test accuracy on clean examples clearly above $90\%$. We use non-targeted $\ell_{\infty}$ norm PGD attack [40] with perturbation norms $\varepsilon_{\infty}=0/255,4/255,8/255,12/255,16/255,24/255,32/255$ to craft adversarial examples (partly visualized in Figures 14(a) and 14(b)). For $\varepsilon_{\infty}=\epsilon/255$, we take $45\epsilon/$ 4 PGD steps with step size $\alpha=2/255$.
Furthermore, recall that in Section 4.3.4, we mentioned that one possible way to trade robustness for clean accuracy is to use a smaller $t<T$. Here, we visualize some examples to further interpret this trade-off, in Figure 14. We can see that using a large defense strength (a large $t$), BCM performs well in eliminating all adversarial noises. However, as the attack budget increases, the purified images start to lose more details. On the other hand, using a small defense strength, BCM cannot remove all adversarial noises when the attack is strong. This explains the clean-robust accuracy trade-off observed in Figure 13.
We additionally remark that on the side of the adversarial defense, such trade-off is also a well-known phenomenon in white-box settings: learning to defend against very strong adversarial attacks significantly increases learning difficulty that may cause underfitting and harm natural accuracy [53, 54, 55]. Moreover, models trained with weak adversarial attacks (e.g., $\varepsilon_{\infty}=4/255$) usually exhibit poor robustness to stronger adversarial attacks (e.g., $\varepsilon_{\infty}=8/255$), so the model must be retrained with stronger adversarial attacks to gain robustness against stronger adversarial attacks, simultaneously with lower clean accuracy. In this sense, going back to the black-box settings, our BCM does not need to be retrained or further sacrifice clean accuracy for robustness against stronger attacks, which might be a worth-exploring advantage.
Appendix B Derivation of the Network Parameterization

### v4

@@ Appendix A Algorithms
Algorithm 7 BCM’s inpainting with refinement
Network $\bm{f}_{\bm{\theta}}(\cdot,\cdot,\cdot)$, BCM inpainting time steps $\varepsilon=t_{1}<\dots<t_{N}\leq T$, Refinement time steps $\tau_{1}>\dots>\tau_{M}$, initial image ${\mathbf{x}}$, binary image mask $\bm{\Omega}$ where 1 indicates the missing pixels, initial noise scale for masked region $s$.
Inpainted image $\hat{{\mathbf{x}}}$.
# Initialize with BCM’s inpainting
$\hat{{\mathbf{x}}}\leftarrow$ Algorithm 6$(\bm{f}_{\bm{\theta}},t_{1},\dots,t_{N},{\mathbf{x}},\bm{\Omega},s)$.
# Refine
for $m=1,\cdots,M$ do
     $\hat{{\mathbf{x}}}\sim\mathcal{N}(\hat{{\mathbf{x}}},\tau_{m}^{2}{\bm{I}})$.
     $\hat{{\mathbf{x}}}\leftarrow\bm{f}_{\bm{\theta}}(\hat{{\mathbf{x}}},\tau_{m},0)$.
     $\hat{{\mathbf{x}}}\leftarrow{\mathbf{x}}\odot(1-\bm{\Omega})+\hat{{\mathbf{x}}}\odot\bm{\Omega}$.
end for
Return: $\hat{{\mathbf{x}}}$.
Appendix B Additional Experiments
B.1 Additional Applications with Bidirectional Consistency
Here, we provide more demonstrations of BCM’s application.
B.1.1 Inpainting on Higher-Resolution Images
(a) FFHQ $64\times 64$
(b) LSUN $256\times 256$
Figure 8: Inpainting on FFHQ and LSUN Bedroom. We compare inpainting with BCM (Algorithm 6 and Algorithm 7) with CM’s inpainting. We visualize more uncurated examples in figs. 24 and 25.
In the main text, we demonstrate that BCM can yield better inpainting results with fewer NFEs on CIFAR-10 images compared with CM’s inpainting algorithm proposed by Song et al. (2023). On high-resolution images, however, we find that simply following the same algorithm as CIFAR-10 (Algorithm 6) is suboptimal, possibly because the correlation between pixels is smaller than that in low-resolution. Fortunately, we find Algorithm 6 can quickly fill the missing region with imperfect but sensible results. Therefore, we can add a few more steps to refine the inpainting results by iteratively adding some small noise to the inpainting regions and denoising by BCM. We summarize this algorithm in Algorithm 7. This can be viewed as a combination of BCM’s inpainting by Algorithm 6 with CM’s inpainting algorithm proposed by Song et al. (2023). We observe that this combination can reduce the NFEs without hurting the inpainting performance.
We evaluate Algorithms 6 and 7 on FFHQ $64\times 64$ and LSUN Bedroom $256\times 256$ in Figure 8 and also compare the results of CM’s inpainting. The hyperparameters are included in Section C.3.2. As we can see: 1) simply adopting BCM’s inpainting as described in Algorithm 6 can fill in relatively sensible content but is imperfect in the overall color and details; 2) CM’s inpainting algorithm yields satisfactory performance with significantly more NFEs; and 3) applying BCM’s inpainting with refinement as described in Algorithm 7 can achieve a comparable performance with fewer NFEs.
B.1.2 Blind restoration of compressed image.
BCM’s bidirectional consistency can offer broader applications. For example, we can invert an image with OOD artifacts to the noise space and re-generate the image from the noise. During inversion, the model maps OOD regions to in-distribution noise, effectively eliminating these artifacts. One potential application leveraging this property is blind restoration of compressed image, where OOD artifacts originate from compression. We show this potential with a proof-of-concept experiment by applying BCM to restore images compressed by JPEG. We present the results in Figure 9, and also include two established restoration approaches, I2SB (Liu et al., 2023) and Palette (Saharia et al., 2022), for comparison. We include hyperparameters for our approach and these baselines in Section C.3.3. Surprisingly, BCM achieves better FID compared to both I2SB and Palette when NFE is limited.
However, since there is no unconditional ADM weight for ImageNet-$64\times 64$ and our BCM is trained with class conditions, we include labels in both the baseline methods and our method to ensure convenience and fair comparison. This inclusion may lead to an overestimation of classifier accuracy or Inception Score (IS). Therefore, we only compare FID.
To capture content changes, we also measure the MSE between restored images and ground truth, motivated by the known rate-perception-distortion trade-off (Blau & Michaeli, 2018; 2019). BCM achieves a per-pixel MSE of about 0.004, while I2SB models achieve about 0.003. With a limited NFE, I2SB further reduces MSE to 0.002, though with highly over-smoothed content. Due to this trade-off, it is difficult to say which model is superior. However, unlike I2SB and Palette, which require paired data with known degradation for training, BCM operates in a fully zero-shot and blind manner. This highlights its potential for downstream inversion tasks.
Here, we note a correction compared to our previous manuscript. In the previous version, we report both models achieving an MSE of about 0.004 and conclude that when NFE is limited, our BCM obtains a better restoration than I2SB in a Pareto sense. We obtained this wrong number by comparing restored images with JPEG images instead of the uncorrupted images. Therefore, we cannot conclude whether BCM or I2SB is superior. However, given that BCM does not require any tuning on the paired data, the main conclusion will still hold.
Figure 9: JPEG (QF20) restoration with I2SB, Palette and BCM on ImageNet-64.
B.2 Additional Ablation Studies and Conclusions
Figure 10: Tracking the loss with different objection functions. We include the loss curve of iCT (Song & Dhariwal, 2024) for reference. We can see that the model with BCT loss defined in Equation 10 converges well. Conversely, the model applying Equation 8 instead of Equation 9 for the soft constraint has a much higher loss at the end of the optimization. While the one without CT loss totally diverges.
B.2.1 Comparison between Loss defined with Equation 8 and Equation 9
(a) original images
(b) Inversion w. NFE=1
(c) Inversion w. NFE=4
(d) Inversion w. NFE=12
We track the loss by models trained with both choices in Figure 10, where we can see that the model trained with Equation 8 features a much higher loss in the end. This echoes its failure in the inversion process: as shown in Figure 11, the model trained with Equation 8 fails to provide an accurate inversion. This is because Equation 8 contains two trajectories, starting from ${\mathbf{x}}_{u}$ and ${\mathbf{x}}_{t}$. While both of them are along the SDE trajectories starting from the same ${\mathbf{x}}_{0}$, they do not necessarily reside on the same PF ODE trajectory; in fact, the probability that they are on the same PF ODE trajectory is 00. On the contrary, Equation 9 bypasses this issue since it only involves trajectories starting from the same ${\mathbf{x}}_{t}$.
B.2.2 Ablation of CT loss
Recall our final loss function has two terms, the soft trajectory constraint term and the CT loss term. We note that the soft constraint defined in Equation 9 can, in principle, cover the entire trajectory, so it should also be able to learn the mapping from any time step $t$ to 00, which is the aim of CT loss. However, we find it crucial to include CT loss in our objective. We provide the loss curve trained without CT loss term in Figure 10, where we can see the training fails to converge. For further verification, we also visualize the images generated by the model trained with full BCM loss and the model trained without CT loss term after 200k iterations in Figure 12. We can clearly see that the model without CT loss cannot deliver meaningful outcomes.
B.2.3 Albation of different sampling schemes
In this section, we compare image generation quality using different sampling schemes we proposed in Section 3.3. Specifically, we compare FID v.s. NFE on CIFAR-10 using ancestral sampling, zigzag sampling and their combination in Figure 13. As we can see, both zigzag sampling and ancestral sampling offer performance gain over single-step generation. However, as NFE increases, zigzag sampling and ancestral sampling cannot provide further improvement in generation quality. This phenomenon has also been discussed by Kim et al. (2024). However, the combination of both can yield better FID than either of them in isolation.
Figure 13: Comparing zigzag sampling, ancestral sampling, and their combination on CIFAR-10. Both zigzag sampling and ancestral sampling offer performance gain over single-step generation, where their combination is better than either in isolation.
B.2.4 Coverage of Trajectory during Training
This also points out some future directions to improve BCM. For example, we can design a better sampling strategy during training to ensure a better coverage of the entire trajectory. Or using different sampling strategy for CT loss term and the soft trajectory constraint term.
Appendix C Experiment Details
C.1 Training Settings
CIFAR-10. For all the experiments on CIFAR-10, following the optimal settings in Song & Dhariwal (2024), we use a batch size of 1,024 with the student EMA decay rate of $0.99993$, scale parameter in Fourier embedding layers of 0.02, and dropout rate of 0.3 for 400,000 iterations with RAdam optimizer Liu et al. (2020) using learning rate $0.0001$. We use the NCSN++ network architecture proposed by Song et al. (2021b), with the modification described in Section 3.1. In our network parameterization, we set $\sigma_{\text{data}}=0.5$ following Karras et al. (2022) and Song et al. (2023). Regarding other training settings, including the scheduler function for $N(\cdot)$, the sampling probability for $t_{n}$ (aka the noise schedule $p(n)$ in (Song et al., 2023; Song & Dhariwal, 2024)), the distance measure $d(\cdot,\cdot)$, we follow exactly Song & Dhariwal (2024), and restate below for completeness’s sake:
$K$ is the total training iterations
FFHQ $64\times 64$ and LSUN Bedroom $256\times 256$. Except the differences stated below, we train BCMs on FFHQ $64\times 64$ and LSUN Bedroom $256\times 256$ using exactly the same settings:
  • For FFHQ $64\times 64$, we use a batch size of 512 but still train the model for 400,000 iterations.
  • For LSUN Bedroom $256\times 256$, we (i) use a dropout rate of 0.2, (ii) adopt the network architecture configuration on this dataset in Song et al. (2021b), (iii) use a batch size of 128, and (iv) adopt a different discretization curriculum from Equation 14. Specifically, in Equation 14, the training procedure is divided into 8 stages. Each stage uses a different $N$ and has $K/8$ iterations. In our experiments, we still divide the entire training procedure into 8 stages, and still set $N(k)=\min(s_{0}2^{j-1},s_{1})+1$ at stage $j$, for $j=1,2,\cdots,8$. However, we do not divide these stages evenly. Instead, we assign 75,000, 60,000, 45,000, 30,000 iterations to the 1st/2nd stage, 3rd/4th stage, 5th/6th stage, 7th/8th stage, respectively  [^3] We assign more iterations to earlier stages with small $N$, as we empirically find that it yields better performance given limited computing budget.. The total number of iterations is 420,000.
Moreover, since our experiments include many applications that are necessarily conducted on test sets (e.g., inpainting), we split FFHQ into a training set of 69,000 images and a test set containing the last 1,000 images. For LSUN Bedroom, we use its official validation set as test set. We implement our model and training algorithm based on the codes released by Song et al. (2023) at https://github.com/openai/consistency_models_cifar10 (Apache-2.0 License).
We specially emphasize that the settings we use to train FFHQ $64\times 64$ and LSUN Bedroom $256\times 256$ are not fine-tuned to optimal (in fact, most of them have not been tuned at all). For instance, Song et al. (2023) trained LSUN Bedroom $256\times 256$ with a significantly larger batch size of 2,048 for 1,000,000 iterations, meaning that their model has seen more than 38×38$\times$38 × samples than ours during training. Therefore, we note that our settings for these two datasets should NOT be used as guidance for training the models to optimal performance, and our results are only for demonstration purposes and should NOT be directly compared with other baselines unless under aligned settings.
ImageNet-64. Instead of training from scratch, we apply bidirectional consistency fine-tuning on ImageNet-64. Specifically, we first train an iCT on ImageNet-64, and then initialize and fine-tune BCM from it. Since there is no official implementation or available checkpoints for iCT, we reproduce it by ourselves. Therefore, we first state our training settings for iCT:
Reproducing iCT on ImageNet-64. Our training setting follows Song & Dhariwal (2024) closely. Most of the training parameters for ImageNet-64 are consistent with those used for training BCM on CIFAR-10, with the exceptions noted below:
  • We use a batch size of 4096 and train the model for 800,000 iterations.
  • The EMA decay rate is set to 0.99997.
  • We adopt the ADM architecture and remove AdaGN following Song & Dhariwal (2024).
  • Instead of Fourier Embeddings, we employ the default positional embedding.
  • The dropout rate is set to 0.2 and only applied to convolutional layers whose feature map resolution is smaller or equal to $16\times 16$.
  • We use mix-precision training.
Additionally, when reproducing iCT on ImageNet-64, following EDM2 (Karras et al., 2024), we slightly modify the network architecture as follows:
  • The self-attention layer at resolution $32\times 32$ is removed.
  • The $Q,K$ and $V$ vectors are normalized in the self-attention layers.
We empirically find these simple modifications bring positive influences on performance.
Fine-tuning BCM on ImageNet-64. We initialize our BCM with the pretrained iCT weights and fine-tune it using Equation 10. Recall that BCM takes in three arguments: 1) the sample ${\mathbf{x}}_{t}$ at time $t$, 2) the current time step $t$, and 3) the target time step $u$, and outputs the sample at time $u$, i.e., ${\mathbf{x}}_{u}$. In contrast, the pretrained iCT model only takes the first two inputs and can be viewed as a special case of BCM when the target time step $u$ is set to 00. Therefore, we aim to initialize BCM such that it preserves the performance of the pretrained CM when $u=0$. To achieve this, we carefully initialize the BCM as follows:
  • In iCT, the time embedding of $t$ is passed through a 2-layer MLP with SiLU activation, after which it is fed into the residual blocks. At the initialization of BCM, we duplicate the 2-layer MLP (with the same pretrained weights) for $u$. However, during fine-tuning, the MLPs for $t$ and $u$ are not tied.
  • We do not want to expand the input dimension of the pretrained residual blocks, so first concatenate the embedding of $t$ and $u$ together, and then reduce its dimension by half with a simple linear layer (without bias). No additional activation function is applied for this simple linear layer.
  • In order to maintain the generation quality of the pretrained iCT, we initialize this linear projection as $[{\mathbf{I}},\bm{0}]$, in which ${\mathbf{I}}$ represents an identity matrix and $\bm{0}$ represents a zero matrix. As a result, the influence brought by the newly introduced time embedding of $u$ is zeroed at the beginning and learned gradually during fine-tuning.
With the modified architecture initialized with the pretrained CM weights, we tune the models using our proposed BCT loss till convergence, using the same learning rate, ema rate, etc., as in iCT. The only difference is that we start the fine-tuning from $N=320$ and increase it to $N=480$ and subsequently $N=640$. Specifically,
  • we tune BCM for 210k, 16k and 8k iterations respectively at $N=320,480$ and 640.
  • we tune BCM-deep for 360k, 100k and 10k iterations respectively at $N=320,480$ and 640.
The number of iterations for different values of $N$ is determined empirically: we increase $N$ when we observe no performance improvements by extending the training duration at the same $N$.
C.2 Sampling and Inversion Configurations
Here we provide hyperparameters for all the experiments on CIFAR-10 and ImageNet-$64\times 64$ to reproduce the results in the main paper.
Sampling. On CIFAR-10: for BCM (not deep), we use ancestral sampling with $t_{1}=1.2$ for NFE$=2$, zigzag sampling with $\varepsilon_{1}=0.2,t_{1}=0.8$ for NFE$=3$ and the combination of ancestral sampling and zigzag sampling with $t_{1}=1.2,\varepsilon_{1}=0.1,\tau_{1}=0.3$ for NFE$=4$. For BCM-deep, we use ancestral sampling with $t_{1}=0.7$ for NFE$=2$, zigzag sampling with $\varepsilon_{1}=0.4,t_{1}=0.8$ for NFE$=3$ and the combination with $t_{1}=0.6,\varepsilon_{1}=0.14,\tau_{1}=0.3$ for NFE$=4$. On ImageNet-$64\times 64$: for both conditional BCM and BCM-deep, we use ancestral sampling with $t_{1}=2.4$ for NFE$=2$, zigzag sampling with $\varepsilon_{1}=0.1,t_{1}=1.2$ for NFE$=3$ and the combination of ancestral sampling and zigzag sampling with $t_{1}=3.0,\varepsilon_{1}=0.12,\tau_{1}=0.4$ for NFE$=4$.
Inversion. On CIFAR-10: we set $\varepsilon=t_{1}=0.07$, $t_{2}=6.0$ and $t_{3}=T=80.0$ for NFE$=2$ and $t_{2}=1.5,t_{3}=4.0,t_{4}=10.0,t_{5}=T=80.0$ for NFE$=4$. These hyperparameters are tuned on 2,000 training samples, and we find them generalize well to all test images. On ImageNet-$64\times 64$: for both BCM and BCM-deep, we set $\varepsilon=t_{1}=0.07$, $t_{2}=15.0$ and $t_{3}=T=80.0$ for NFE$=2$. In the experiments we always use 1-step generation to map the inverted noise to reconstructed images (though using more than one step may improve the results) and evaluate the per-dimension MSE between the original images and their reconstructed counterparts.
We highlight that the hyperparameters are relatively robust for 2-step sampling/inversion, and the trend that the combination of ancestral and zigzag sampling is superior is also general. However, to achieve optimal performance with more steps, the tuning of each specific time step may require great effort. We should note that similar effort is also required in CMs, where Song et al. (2023) use ternary search to optimize the time steps.
Inversion Baselines. For DDIM, we use the reported MSE by Song et al. (2021a); for EDM, we load the checkpoint provided in the official implementation at https://github.com/NVlabs/edm?tab=readme-ov-file (CC BY-NC-SA 4.0 License), and re-implement a deterministic ODE solver following Algorithm 1 in (Karras et al., 2022).
C.3 Applications with Bidirectional Consistency
Here we provide more details about the applications we demonstrated in Section 4.3 and Section B.1.
C.3.1 Interpolation
We first invert the two given images ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$ to noise at $T=80.0$ using Algorithm 5. To avoid subscript overloading, in this section, we denote their noise as ${\mathbf{z}}_{1}$ and ${\mathbf{z}}_{2}$, respectively. Specifically, we find that adopting a 3-step inversion with $\varepsilon=t_{1}=0.07,t_{2}=1.5,t_{3}=6.0$, and $t_{4}=T=80.0$ for CIFAR-10, and a 2-step inversion with $\varepsilon=t_{1}=0.07,t_{2}=1.8$, and $t_{3}=T=80.0$ for FFHQ and LSUN bedroom are sufficient for good reconstruction results. Note, that we use two different small initial noise, i.e., $\bm{\sigma}_{1},\bm{\sigma}_{2}\sim\mathcal{N}(0,{\bm{I}}),\bm{\sigma}_{1}\neq\bm{\sigma}_{2}$ when inverting ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$ respectively.
Since BCM learns to amplify the two initial Gaussian i.i.d. noises, it is reasonable to hypothesize that the amplified noises (i.e., the embeddings) ${\mathbf{z}}_{1}$ and ${\mathbf{z}}_{2}$ reside on the same hyperspherical surface as if ${\mathbf{z}}_{1}$ and ${\mathbf{z}}_{2}$ are directly sampled from $\mathcal{N}(0,T^{2}{\bm{I}})$. Therefore, following Song et al. (2023), we use spherical linear interpolation as
As we discussed in the main text, it is crucial to set $\bm{\sigma}_{1}\neq\bm{\sigma}_{2}$ for inversion. A possible reason is that if using $\bm{\sigma}_{1}=\bm{\sigma}_{2}$ for inversion, the inverted noises ${\mathbf{z}}_{1}$ and ${\mathbf{z}}_{2}$ may reside on an unknown submanifold instead of the hyperspherical surface of Gaussian, and hence Equation 16 cannot yield ideal interpolation results. In Appendix F, we present visualization and discussions on the geometric properties of the noise space.
C.3.2 Inpainting
@@
We describe BCM’s inpainting process in Algorithm 6. In our experiments on CIFAR-10, we set the initial noise scale $s=0.5$. Additionally, instead of inverting the image back to $T=80.0$, we empirically find inverting the image to $T=2.0$ already suffice for satisfactory inpainting outcomes, and hence, we use a 3-step inversion (i.e., $N=4$), where $t_{1}=\varepsilon=0.07$, $t_{2}=0.4$, $t_{3}=1.0$ and $t_{4}=T=2.0$, and 1-step generation from $T$ to 00.
@@
As for the results by CMs, we implement CM’s inpainting algorithm (Algorithm 4 in Song et al. (2023)). The total number of time steps (NFEs) is set to 18 according to the official inpainting script at https://github.com/openai/consistency_models_cifar10/blob/main/editing_multistep_sampling.ipynb. To ensure a fair comparison, we opt for the improved model, iCT (Song & Dhariwal, 2024), over the original CMs, since iCT delivers superior generation performance. Nonetheless, due to the absence of officially released codes and checkpoints by the authors, we reproduce and train our own iCT model. Our reproduced iCT yields an FID score of single-step generation closely matching that reported by Song & Dhariwal (2024) — our reproduced FID is $2.87$ compared to the reported $2.83$ — affirming the reliability of our results.
For higher-resolution images, we summarize BCM’s inpainting with refinement in Algorithm 7. For a fair comparison among BCM’s inpainting algorithm, CM’s inpainting algorithm and BCM’s inpainting with refinement, we run these three algorithms on the same BCM trained on FFHQ $64\times 64$/LSUN Bedroom $256\times 256$. Then, we perform the inpainting using the following hyper-parameters:
  • FFHQ $64\times 64$: BCM inpainting: we adopt initial noise scale $s=1.0$, 1-step inversion with $t_{1}=\varepsilon=0.5$ and $t_{2}=T=20.0$, and 1-step generation from $T$ to 00 in BCM’s inpainting. CM inpainting: we adopt a 24-step noise schedule following Song et al. (2023). BCM inpainting w. refine: we first perform BCM inpainting, followed by refinement using only the last $M=12$ time steps used in CM inpainting.
  • LSUN Bedroom $256\times 256$: BCM inpainting: we use initial noise scale $s=0.5$, 2-step inversion with $t_{1}=\varepsilon=0.5,t_{2}=2.5$ and $t_{3}=T=200.0$, and 2-step ancestral generation with $t_{1}=4.4$ in BCM’s inpainting. CM inpainting: we adopt a 40-step noise schedule following Song et al. (2023). we adopt the same 40-step noise schedule in Song et al. (2023). BCM inpainting w. refine: we first perform BCM inpainting, followed by refinement using only the last $M=26$ time steps used in CM inpainting.
C.3.3 JPEG Restoration
We invert the input image to reach the noise scale of $T=3.2$ using $1/2/3$ steps and reconstruct it by single-step generation, making a total NFE of $2/3/4$ as in Figure 9. For 1-step inversion, we use $\varepsilon=t_{1}=0.5,t_{2}=T=3.2$. For 2-step inversion, we use $\varepsilon=t_{1}=0.5,t_{2}=1.2,t_{3}=T=3.2$. For 3-step inversion, we use $\varepsilon=t_{1}=0.5,t_{2}=1.2,t_{3}=2.0,t_{4}=T=3.2$.
For the I2SB (Liu et al., 2023) baseline, since there’s no pretrained checkpoints available for ImageNet-$64\times 64$, we reproduce it using the official codebase at https://github.com/NVlabs/I2SB (NVIDIA Source Code License). There is also no official implemention or checkpoints available for Palette (Saharia et al., 2022), so we also reproduce it by modifying from the I2SB codebase. We initialize both methods with a pretrained ADM model with weight available at https://github.com/openai/guided-diffusion (MIT License).
Appendix D Derivation of the Network Parameterization